Two cards drawn from standar deck of cards.Find probability
P(1 heart, 1club)
P(1 queen, 1 king)
Thank you so much

Answer by ewatrrr(24785)


Hi
Two cards drawn from standard deck of cards.
Note: will assume this is done 'without replacing' first card.
P(1 heart, 1club) = 13/52*13/51
P(1 queen, 1 king) = 4/52*4/51
If the first card were replaced prior to drawing second card...
P(1 heart, 1club) = 13/52*13/52
P(1 queen, 1 king) = 4/52*4/52
Always important to note if experiement is done with or w/o replacement